OpenStudy (anonymous):

Find the thirteenth term of an arithmetic sequence if the first term is 3 and the common difference is 3

OpenStudy (anonymous):

Arithmetic sequence is addition of 3 So the sequence is 3,6,9,12,15,18,21,24,27,30,33,36,39 The thirteenth term of this arithmetic sequence is 39

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{13} = 3 + (13 - 1)3\] \[a _{13} = 3 + (12)3\] \[a _{13} = 3 + \] \[a _{13} = 39\]

OpenStudy (anonymous):

Thank you very much.

OpenStudy (anonymous):

@omgdime yes you were correct, but your way is not the best way. Since they can ask you for the 150 in an exam and you wouldn't want to spend too much time doing it your way.

OpenStudy (anonymous):

I know, I was just trying to help her. not so much as work it out. I should have though @some_someone

OpenStudy (anonymous):

lol @omgdime i know you were just helping ;D & its good.

OpenStudy (anonymous):

lol thank you @some_someone

OpenStudy (anonymous):

@omgdime , thank you for your help :)

OpenStudy (anonymous):

you welcome sweethear :)